# Optical and UV Flares from Binary Black Hole Mergers in Active Galactic Nuclei

### v1

@@ III Emission profiles
Setting $f_{\mathrm{c}}=3$, we obtain an AGN spectrum consistent with the emission of the corresponding SS thin disc (see Figures 4-5).
@@
Non-blazar AGNs typically exhibit stochastic variability in optical and X-ray bands representing about 10% of their average emission over time scales ranging from days to years [42, 43]. The origin of such variability is not established, as it could be driven by multiple processes in the AGN disc. Therefore, here we consider $r_{\mathrm{th}}=0.25$ as a reasonable threshold to distinguish the flares predicted in this work from their host emission, and define as detectable emission solutions leading to $r_{\nu}\geq r_{\mathrm{th}}$ (see equation 38).
In the following subsections, we illustrate spectral energy distributions (SEDs) and LC profiles predicted by the present model using a fiducial BH remnant of $40$ M with inflow/outflow efficiencies of $\eta_{\mathrm{in}}=0.1$ and $\eta_{\mathrm{w}}=0.005$. The other model parameters are varied as indicated.
@@ III.1 Spectral energy distributions
The EM counterpart discussed here is approached as blackbody emission from the photosphere of the disc eruption (see equation 34). As such, the appropriate frequency for observing these flares is $\nu_{\mathrm{pk}}\sim k_{\mathrm{B}}T_{\mathrm{eff}}/[(1+z)h]$ and somewhat lower frequencies. We note that, at maximum bolometric luminosity, the effective temperature $T_{\mathrm{eff}}$ scales inversely with the disc height (see equations 29-33). Therefore, we expect the flares to shine at lower frequencies as the hosting SMBH is larger. This is illustrated in Figures 4-5, where we present SEDs of the disc eruptions considering AGNs with central engines of $10^{6}$ and $10^{7}$ M.
The SEDs displayed in Figure 4 correspond to an “upper eruption” of a merger with a kick angle of $\theta_{\mathrm{k}}=5^{\circ}$, in an AGN with $M_{S}=10^{6}$ M, located at $D_{\mathrm{L}}=1000$ Mpc. In each frame, we display the SEDs of the disc eruption at different observed times after the detection of the GW event, as labelled, starting with the time when photons can escape from the plasma ejection (given by equation 35). We over-plot the spectra of the associated disc emission (obtained with the SS model, thick dashed curve) and of the host galaxy given by equation 39 (thick solid curve). The vertical lines indicate optical ($g$-band), near UV, and far UV frequencies, as labelled. Left/right panels of Figure 4 consider different merger locations $a$ and upper/lower panels different kick velocities $v_{\mathrm{k}}$, as indicated. These values for $a$ and $v_{\mathrm{k}}$ are chosen so that condition 37 is fulfilled.
@@
Figure 5 shows SEDs of the merger flare with the same parameter set as in Figure 4, but considering a central SMBH of $10^{7}$ M, and merger locations and kick velocities consistent with condition 37. As expected, these flares peak at lower frequencies than in the $M_{\mathrm{S}}=10^{6}$ M case, and can be suitably detected at optical and UV frequencies. On the other hand, the time scale for the flare lags and durations increases by about an order of magnitude with respect to the $M_{\mathrm{S}}=10^{6}$ M case.
Following the trend exhibited by the SEDs of Figures 4 and 5, one might expect that mergers in AGNs with SMBHs of $10^{8}$ M and heavier can be detectable at optical and infrared frequencies. Nevertheless, the flare time lags in these larger AGNs would be of year scales (we analyse time lag and time duration in more detail in Section III.3), and therefore the multi-messenger associations of real data with such characteristics could be non-trivial. We therefore suggest that the emission mechanism discussed in this paper is suitable for seeking EM counterparts in hosts with $M_{\mathrm{S}}\sim 10^{6-7}$ M.
Figure 4: SED snapshots of the disc eruption (the merger EM counterpart) in an AGN with a central engine of $10^{6}$ M. The SEDs are calculated at different post-merger time steps $t_{\mathrm{pGW}}$ as indicated by the legends in each frame. The emission of the associated thin disc and host galaxy are over-plotted with the thick dashed and solid curves, respectively. The vertical lines indicate the optical g-band, near UV, and far UV frequencies. Curves on the left and right panels are calculated assuming a merger at $9000$ and $20000R_{\mathrm{g}}$ from the central SMBH, whereas upper and lower panels consider kick velocities of 300 and 500 km s-1. Other model parameters shared by all the panels are specified in the text.
@@ III.2 Optical light-curves
Figure 6 displays the profiles of possible LCs driven by a BBH merger at $5000$ $R_{\mathrm{g}}$ from an SMBH of $10^{6}$ M with a kick angle of $\theta_{\mathrm{k}}=5^{\circ}$ of the merger remnant. Curves in the left and right panels of Figure 6 are obtained assuming an AGN accreting at 0.01 and 0.05 Eddington accretion rates, respectively. In each frame, upper, middle, and lower curve families are obtained assuming different redshifts of the source, as indicated. Curves with different line styles are obtained using different kick velocities as labelled, whereas red and blue curves correspond to the upper and lower disc eruptions. The LCs are calculated as given by equation 41 and converted to AB magnitudes employing the Python package Speclite [^2] https://speclite.readthedocs.io/en/latest/index.html with the $g$ filter of the Sloan Digital Sky Survey (SDSS, Blanton et al. 80). For reference, the horizontal lines on the right in each panel indicate the limiting magnitude of different optical instruments, namely DDOTI (1000s exposure, 3-$\sigma$, white filter [10, 81]), ZTF (30s exposure, 5-$\sigma$ detection, g filter [39]), SPLUS (3$\times$30s, 3-$\sigma$, g filter [82, 83]), DECam (60s exposure, 5-$\sigma$ detection, g filter [28]), and the forthcoming Vera C. Rubin Observatory ($2\times 15$s exposure, 5-$\sigma$ detection, g filter [84]), as labelled.
@@
Figure 6 shows how the flare profiles are impacted by the remnant’s kick velocity $v_{\mathrm{k}}$ and by the SMBH accretion rate $\dot{m}_{\mathrm{S}}$. The emission time lag is shorter for larger values of $v_{\mathrm{k}}$ since the disc crossing time (and hence the time for the production of the disc eruptions) is reduced (see equation 26). Furthermore, the amplitude of the flare is reduced as $v_{\mathrm{k}}$ increases since the energy stored in the disc ejections is approximately proportional to $v_{\mathrm{k}}^{-4}$ (see equation 10). We note that the flux enhancement relative to the (here assumed) stationary AGN emission is smaller as the AGN accretion rate $\dot{m}_{\mathrm{S}}$ is larger. This suggests that EM counterparts can be more distinguishable from their hosts in discs with relatively low accretion rates. On the other hand, for larger $\dot{m}_{\mathrm{S}}$ the overall flux (i.e., AGN plus merger counterpart) increases, hence increasing the distance up to which the source can be detected. These LCs show, for instance, that in AGNs of $10^{6}$ M accreting at $\dot{m}_{\mathrm{S}}=0.01-0.05$, the flares could be detected by ZTF up to $z\sim 0.04-0.09$, and up to $z\sim 0.1-0.3$ by forthcoming Rubin. The curves show that upper eruptions appear always somewhat before and have shorter durations than the lower counterparts. This is because upper ejections are always less massive than the lower ones, and therefore their photon diffusion time (see equations 30 and 27) and flare characteristic period (equation 36) are always shorter in the upper flares.
@@
In Figure 7, we present possible LCs of a flare triggered by a merger with the same parameters as in Figure 6, but considering an SMBH of $10^{7}$ M, and exploring different values for the kick velocity and the redshift of the source, as indicated. The flare time lags and durations are of the order of hundreds and tens of days, respectively, which represents an increase of about an order of magnitude compared to such periods in Figure 6. However, the effect on the LC profiles due to variations of $v_{\mathrm{k}}$ and $\dot{m}_{\mathrm{S}}$ is qualitatively similar. We also note that in this $M_{\mathrm{S}}=10^{7}$ M case, lower values for $v_{\mathrm{k}}$ are allowed by condition 37, and that the overall emission can be detected at larger redshifts compared to the $M_{\mathrm{S}}=10^{6}$ M case. For instance, curves in Figure 7 show that in discs accreting at $\dot{m}_{\mathrm{S}}=0.01-0.05$, the flares could be detected by ZTF up to $z\sim 0.1-0.3$, and up to $z\sim 0.5-1$ by forthcoming Rubin.
@@
The SEDs and LCs presented in Figures 4-7 suggest that, in general, the flares are more distinguishable from their host as the merger distance $a$ and the kick velocity $v_{\mathrm{k}}$ are smaller.
@@
Figure 6: LCs in the optical $g$-band of the EM counterpart model for a BBH merger in an AGN disc. The figure illustrates possible LC profiles driven by a merger with the following model parameters: $M_{\mathrm{S}}=10^{6}$ M, $a=5000R_{\mathrm{g}}$, $M_{\bullet}=40$ M, $\theta_{\mathrm{k}}=5^{\circ}$, $\eta_{\mathrm{in}}=0.1$ and $\eta_{\mathrm{w}}=0.005$ (see the text for details). We explore different redshifts for the source which results in the upper, middle, and lower curve families as well as different kick velocities, as labelled. Red and blue curves correspond to the “upper” and “lower” eruptions of the model. Left and right panels correspond to AGNs accreting at different rates $\dot{m}_{\mathrm{S}}$, as indicated. The horizontal lines indicate limiting magnitudes of different optical instruments as labelled (see details in the text).
@@
Figure 7: Same as in Figure 7, but considering an SMBH of $M_{\mathrm{S}}=10^{7}$ M and different kick velocities $v_{\mathrm{k}}$ and source redshifts $z$, as indicated.
@@
In Figure 8, we display $\Delta t_{\ell}$ given by equation 35 as a function of the SMBH mass $M_{\mathrm{S}}$. In this example, we consider a source located at $D_{\mathrm{L}}=1000$ Mpc and derive the $\Delta t_{\ell}$ vs $M_{\mathrm{S}}$ curves exploring different values for $a/R_{\mathrm{g}}$, $v_{\mathrm{k}}$, $\theta_{\mathrm{k}}$, and $\dot{m}_{\mathrm{S}}$. We restrict these curves to fluxes with ratios $r_{\nu}\geq 0.25$ in the optical $g$-band (see equation 41) together with condition 37. Thus, the curves in Figure 8 initiate/terminate when the aforementioned conditions are no longer satisfied. In each panel, the curves with different values of $a/R_{\mathrm{g}}$ are labelled with different line styles, whereas curves of different kick velocities are labelled with different colours. Computations in the upper and lower panels are obtained assuming different kick angles $\theta_{\mathrm{k}}$, whereas left and right panels correspond to different SMBH accretion rates $\dot{m}_{\mathrm{S}}$, as labelled.
@@
The curves in Figure 8 show that the time lag of the EM counterpart is approximately correlated to the SMBH mass as $\propto M_{\mathrm{S}}^{\alpha_{\ell}}$ with $\alpha_{\ell}\approx 1$. As already noted in the SEDs and LCs of previous subsections, Figure 8 shows that $\Delta t_{\ell}$ increases for larger merger locations and is diminished for larger kick velocities. The parameter space leading to visible flares is reduced for larger values of $\dot{m}_{\mathrm{S}}$ and $\theta_{\mathrm{k}}$, as also noted in the previous subsections. Interestingly, the condition $r_{\nu}\geq r_{\mathrm{th}}$ together with condition 37 associates the mass of the hosting SMBH with a characteristic kick velocity that would produce detectable flares. For instance, we see that in AGNs with SMBHs of $10^{6}$ M, visible flares would be more likely produced by remnants with kicks of $\sim 350$ km s-1, and of $\sim 150$ km s-1 in discs around SMBHs of $10^{7}$ M. Such $M_{\mathrm{S}}-v_{\mathrm{k}}$ association appears more defined as the SMBH accretion rate is larger.
@@
In Figure 9, we display the flare duration given by equation 36 and associated to the curves of Figure 8. This figure shows that $\Delta t_{\mathrm{d}}$ is generally one order of magnitude smaller than the flare time lag. The $\Delta t_{\mathrm{d}}-M_{\mathrm{S}}$ relation is approximately linear, similar to $\Delta t_{\ell}$. However, we note that the effect of the merger location $a$ is less pronounced in the flare duration than in the flare time lag.
@@
The present EM counterpart model is built such that the flare time lag is composed of a sequence of periods, namely the cavity crossing time $\Delta t_{\mathrm{cav}}$, the time for the outflow formation $\Delta t_{\mathrm{HL}}$, the mean time for the formation of the disc ejection $\Delta t_{\mathrm{h}}$, and the photon escaping time $\Delta t_{\mathrm{ph}}$ (see Section II.2 for details). The contribution of these periods to the flare time lag varies according to the model parameter configuration. Generally, $\Delta t_{\mathrm{cav}}$ and $\Delta t_{\mathrm{h}}$ are the dominant components, as illustrated in Figure 10. In this figure, we display the components of the time lag as a function of merger location, considering an AGN at a distance of 1000 Mpc, accreting at a rate of $\dot{m}_{\mathrm{S}}=0.01$, and within the conditions 37 and $r_{\nu}\geq r_{\mathrm{th}}$ (similarly as in Figures 8-9). Curves in the upper, middle, and lower panels are calculated with different kick angles as indicated. Based on the $M_{\mathrm{S}}-v_{\mathrm{k}}$ relation discussed in the previous subsection, the curves in the left panels are obtained assuming an SMBH of $M_{\mathrm{S}}=10^{6}$ M with a remnant of $v_{\mathrm{k}}=350$ km s-1, whereas curves in the right panels assume $M_{\mathrm{S}}=10^{7}$ M and $v_{\mathrm{k}}=150$ km s-1. Figure 10 shows that the period $\Delta t_{\mathrm{h}}$ dominates in the case of very small kick angles. On the other hand, the contribution of $\Delta t_{\mathrm{cav}}$ rapidly increases as $\theta_{\mathrm{k}}$ is larger, becoming dominant for $\theta_{\mathrm{k}}\gtrsim 8^{\circ}$ (still in the small kick angle regime). The photon escape time $\Delta t_{\mathrm{ph}}$ is generally subdominant within the parameter space of interest.
@@
Figure 8: Time lag of the counterpart flare after the GW event as a function of the hosting SMBH mass. Curves in the left and right panels are obtained considering different SMBH accretion rates, whereas upper and lower panels depict different kick angles for the merger remnant, as labelled. In each panel, different curves correspond to different locations of the merger and different kick velocities of the remnant, as indicated. The curves terminate when conditions 37 and $F_{\nu}/F_{\nu,\mathrm{host}}\leq 0.25$ are no longer satisfied (see the text for further details).
Figure 9: Same as in Figure 8, but displaying the $\Delta t_{\mathrm{d}}$ vs $M_{\mathrm{S}}$ curves.

### v2

@@ III Emission profiles
Setting $f_{\mathrm{c}}=3$, we obtain an AGN spectrum consistent with the emission of the corresponding thin disc (see Figures 4-5).
@@
Non-blazar AGNs typically exhibit stochastic variability in optical and X-ray bands representing about 10% of their average emission over time scales ranging from days to years [46, 47]. The origin of such variability is not established, as it could be driven by multiple processes in the AGN disc. Therefore, here we consider $r_{\mathrm{th}}=0.25$ as a reasonable threshold to distinguish the flares predicted in this work from their host emission, and define as detectable, those solutions leading to $r_{\nu}\geq r_{\mathrm{th}}$ (see equation 38).
The emission scenario presented in this paper suggests, in principle, that the largest ratios $r_{\nu}$ would be produced by high-mass BH remnants with low kick velocities, as the flare energy scales as a function of $v_{\mathrm{k}}^{-4}$ and $M_{\bullet}^{2}$ (see equation 15). Nevertheless, towards this regime (low $v_{\mathrm{k}}$ and high $M_{\bullet}$), the condition 37 also becomes increasingly difficult to satisfy, making the emission mechanism discussed here uncertain.
On the other hand, a large $M_{\bullet}$ and low $v_{\mathrm{k}}$ could satisfy the condition 37 if the merger occurs at sufficiently large radial distances $a$ within the disc, as the disc height increases with the distance to the SMBH. However, in such cases, the flare time delay also increases, making the multi-messenger association more challenging. A more detailed analysis of flare delays for detectable signatures is presented in Sections II.2 and IV.
In the following subsections, we illustrate spectral energy distributions (SEDs) and LC profiles predicted by the present model using a fiducial BH remnant of 20 M with inflow/outflow efficiencies of $\eta_{\mathrm{in}}=0.1$ and $\eta_{\mathrm{w}}=$0.05, and wind opening angle of $\theta_{\mathrm{w}}=45^{\circ}$. The other model parameters are varied as indicated.
@@ III.1 Spectral energy distributions
The EM counterpart discussed here is approached as blackbody emission from the photosphere of the disc eruption (see equation 34). As such, the appropriate frequency for observing these flares is $\nu_{\mathrm{pk}}\sim k_{\mathrm{B}}T_{\mathrm{eff}}/[(1+z)h]$ and somewhat lower frequencies. We note that, at maximum bolometric luminosity, the effective temperature $T_{\mathrm{eff}}$ scales inversely with the disc height $h_{\mathrm{d}}$ (see equations 29-33). Therefore, we expect the flares to shine at lower frequencies as the hosting SMBH is larger. This is illustrated in Figures 4-5, where we present SEDs of the disc eruptions considering AGNs with central engines of $10^{6}$ and $10^{7}$ M.
The SEDs displayed in Figure 4 correspond to an “upper eruption” of a merger with a kick angle of $\theta_{\mathrm{k}}=10^{\circ}$, in an AGN with $M_{\mathrm{S}}=10^{6}$ M, $\dot{m}$=0.025, located at $D_{\mathrm{L}}=1600$ Mpc. In each frame, we display the SEDs of the disc eruption at different observed times after the detection of the GW event, as labelled, starting with the time when photons can escape from the plasma ejection (given by equation 35). We over-plot the spectra of the associated thin disc emission (derived from the SG03 model, thick dashed curve) and of the host galaxy given by equation 39 (thick solid curve). The vertical lines indicate optical ($g$-band), near UV, and far UV frequencies, as labelled. Left/right panels of Figure 4 consider different merger locations $a$ and upper/lower panels different kick velocities $v_{\mathrm{k}}$, as indicated. These values for $a$ and $v_{\mathrm{k}}$ are chosen so that conditions 4 and 7 are fulfilled.
@@
Figure 5 shows SEDs of the merger flare with the same parameter set as in Figure 4, but considering a central SMBH of $10^{7}$ M, and merger locations and kick velocities consistent with condition 37. As expected, these flares peak at lower frequencies than in the $M_{\mathrm{S}}=10^{6}$ M case, and can be suitably detected at optical and UV frequencies. On the other hand, the time scale for the flare lags increases by nearly three times with respect to the $M_{\mathrm{S}}=10^{6}$ M case.
Following the trend exhibited by the SEDs of Figures 4 and 5, one might expect that mergers in AGNs with SMBHs of $10^{8}$ M and heavier can be detectable at optical and infrared frequencies. Nevertheless, the flare time lags in these larger AGNs would likely increase to several tens to hundred day scales, which would challenge potential MM associations. We analyse the flare time lag and duration in more detail in Section III.3.
Figure 4: SED snapshots of the disc eruption (the merger EM counterpart) in an AGN with a central engine of $10^{6}$ M. The SEDs are calculated at different times, as indicated by the legends in each frame. The first curve of this family corresponds to the maximum bolometric luminosity of the flare. The emission of the associated thin disc and host galaxy are over-plotted with the thick dashed and solid curves, respectively. The vertical lines indicate the optical g-band, near UV, and far UV frequencies. Curves on the left and right panels are calculated assuming a merger at $9000$ and 18000 $R_{\mathrm{g}}$ from the central SMBH, whereas upper and lower panels consider kick velocities of 300 and 400 km s-1. Other model parameters shared by all the panels are specified in the text.
@@ III.2 Optical light-curves
Figure 6 displays the profiles of possible LCs driven by a BBH merger at $9000$ $R_{\mathrm{g}}$ from an SMBH of $10^{6}$ M with a kick angle of $\theta_{\mathrm{k}}=10^{\circ}$ of the merger remnant. Curves in the left and right panels of Figure 6 are obtained assuming an AGN accreting at 0.025 and 0.1 Eddington accretion rates, respectively. In each frame, upper, middle, and lower curve families are obtained assuming different redshifts of the source, as indicated. Curves with different line styles are obtained using different kick velocities as labelled, whereas red and blue curves correspond to the upper and lower disc eruptions, respectively. The LCs are calculated as given by equation 41 and converted to AB magnitudes employing the Python package Speclite [^2] https://speclite.readthedocs.io/en/latest/index.html with the $g$ filter of the Sloan Digital Sky Survey (SDSS, [88]). For reference, the horizontal lines on the right in each panel indicate the limiting magnitude of different optical instruments, namely DDOTI (1000s exposure, 3-$\sigma$, white filter [10, 89]), ZTF (30s exposure, 5-$\sigma$ detection, g filter [43]), SPLUS (3$\times$30s, 3-$\sigma$, g filter [90, 91]), DECam (60s exposure, 5-$\sigma$ detection, g filter [29]), and the forthcoming Vera C. Rubin Observatory ($2\times 15$s exposure, 5-$\sigma$ detection, g filter [92]), as labelled.
@@
Figure 6 shows how the flare profiles are impacted by the remnant’s kick velocity $v_{\mathrm{k}}$ and by the SMBH accretion rate $\dot{m}_{\mathrm{S}}$. The emission time lag is shorter for larger values of $v_{\mathrm{k}}$ since the disc crossing time (and hence the time for the production of the disc eruptions) is reduced (see equation 25). Furthermore, the amplitude of the flare is reduced as $v_{\mathrm{k}}$ increases since the energy stored in the disc ejections is approximately proportional to $v_{\mathrm{k}}^{-4}$ (see equation 13). We observe that for larger $\dot{m}_{\mathrm{S}}$ the overall flux (i.e., AGN plus merger counterpart) increases, hence increasing the distance up to which the source can be detected. These LCs show, for instance, that in AGNs of $10^{6}$ M accreting at $\dot{m}_{\mathrm{S}}=$0.025 - 0.1, the flares could be detected by ZTF up to $z\sim$ 0.04- 0.1, and up to $z\sim$ 0.25-0.5 by forthcoming Rubin. The curves show that upper eruptions appear always somewhat before and have shorter durations than the lower counterparts. This is because upper ejections are always less massive than the lower ones, and therefore their photon diffusion time (see equations 30 and 27) and flare characteristic period (equation 36) are always shorter in the upper flares.
@@
In Figure 7, we present possible LCs of a flare triggered by a merger with the same parameters as in Figure 6, but considering an SMBH of $10^{7}$ M, and exploring different values for the kick velocity and the redshift of the source, as indicated. The flare time lags falls within 40-20 days, which represents an increase of about 3-5 times compared to such period in Figure 6. However, the effect on the LC profiles due to variations of $v_{\mathrm{k}}$ and $\dot{m}_{\mathrm{S}}$ is qualitatively similar. We also note that in this $M_{\mathrm{S}}=10^{7}$ M case, lower values for $v_{\mathrm{k}}$ are allowed by condition 37, and that the overall emission can be detected at larger redshifts compared to the $M_{\mathrm{S}}=10^{6}$ M case. For instance, curves in Figure 7 show that in discs accreting at $\dot{m}_{\mathrm{S}}=0.025-0.1$, the flares could be detected by ZTF up to $z\sim 0.1-0.3$, and up to $z\sim 0.7-1.2$ by forthcoming Rubin.
@@
The SEDs and LCs presented in Figures 4-7 suggest that, in general, the flares are more distinguishable from their host as the merger location $a$ and the kick velocity $v_{\mathrm{k}}$ are smaller.
@@
Figure 6: LCs in the optical $g$-band of the EM counterpart model for a BBH merger in an AGN disc. The figure illustrates possible LC profiles driven by a merger with the following model parameters: $M_{\mathrm{S}}=10^{6}$ M, $a=9000R_{\mathrm{g}}$, $M_{\bullet}=20$ M, $\theta_{\mathrm{k}}=10^{\circ}$, $\eta_{\mathrm{in}}=0.1$ and $\eta_{\mathrm{w}}=0.05$ (see the text for details). We explore different redshifts for the source which results in the upper, middle, and lower curve families as well as different kick velocities, as labelled. Red and blue curves correspond to the “upper” and “lower” eruptions of the model. Left and right panels correspond to AGNs accreting at different rates $\dot{m}_{\mathrm{S}}$, as indicated. The horizontal lines indicate limiting magnitudes of different optical instruments as labelled (see details in the text).
@@
Figure 7: Same as in Figure 7, but considering an SMBH of $M_{\mathrm{S}}=10^{7}$ M, $a=4000R_{\mathrm{g}}$ and different kick velocities $v_{\mathrm{k}}$ and source redshifts $z$, as indicated.
@@
In Figure 8, we display $\Delta t_{\ell}$ given by equation 35 as a function of the merger location $a$ in the AGN disc. In this example, we consider a source located at $D_{\mathrm{L}}=$1600 Mpc and derive the $\Delta t_{\ell}$ vs $a/R_{\mathrm{g}}$ curves exploring different values for $v_{\mathrm{k}}$, $\theta_{\mathrm{k}}$, $M_{\mathrm{S}}$, and $\dot{m}_{\mathrm{S}}$. We restrict these curves to fluxes with ratios $r_{\nu}\geq 0.25$ in the optical $g$-band (see equation 41) together with condition 37. Thus, the curves in Figure 8 initiate/terminate when the aforementioned conditions are no longer satisfied. In each panel, the curves corresponding to different SMBH masses are labelled with different line styles, whereas curves of different kick velocities are labelled with different colours. Computations in the upper and lower panels are obtained assuming different kick angles $\theta_{\mathrm{k}}$, whereas left and right panels correspond to different accretion rates $\dot{m}_{\mathrm{S}}$, as labelled.
@@
The curves in Figure 8 show that the time lag of the EM counterpart increases monotonically with the merger location as $\propto(a/R_{\mathrm{g}})^{\alpha_{\ell}}$ with $\alpha_{\ell}\approx 1$. As already exhibited by the SEDs and LCs of previous subsections, Figure 8 shows that $\Delta t_{\ell}$ increases as the SMBH mass is larger (at a given $a/R_{\mathrm{g}}$) and diminishes for larger kick velocities. Interestingly, the condition $r_{\nu}\geq r_{\mathrm{th}}$ together with condition 37 associates the mass of the hosting SMBH to a characteristic range of $a/R_{\mathrm{g}}$ within which visible flares could be produced. For instance, we see that in AGNs with SMBHs of $10^{6}$ M, visible flares would be more likely produced at $a\sim 10^{4-5}R_{\mathrm{g}}$, whereas they would be better produced at $a\sim 5\times 10^{2-3}R_{\mathrm{g}}$ for SMBHs of $10^{8}$ M.
@@
In Figure 9, we display the flare duration given by equation 36 and associated to the curves of Figure 8. This figure shows that $\Delta t_{\mathrm{d}}$ is generally one order of magnitude smaller than the flare time lag. We observe that flare duration $\Delta t_{\mathrm{d}}$ increases with the magnitude of the kick velocity $v_{\mathrm{k}}$, different to $\Delta t_{\ell}$.
The present EM counterpart model is built such that the flare time lag is composed of a sequence of periods, namely the cavity crossing time $\Delta t_{\mathrm{cav}}$, the time for the wind ignition $\Delta t_{\mathrm{HL}}$, the time for the wind to expel the disc matter $\Delta t_{\mathrm{bo}}$, and the time for photons to escape the matter ejections $\Delta t_{\mathrm{ph}}$ (see Section II.2 for details). The contribution of these periods to the flare time lag vary according to the model parameter configuration.
@@
In Figure 10, we illustrate the components of the time lag as a function of merger location, considering an AGN at a distance of 1600 Mpc, accreting at a rate of $\dot{m}_{\mathrm{S}}=0.025$, and within the conditions 37 and $r_{\nu}\geq r_{\mathrm{th}}$ (similarly as in Figures 8-9). Curves in the upper, middle, and lower panels are calculated with different kick angles as indicated. Based on the $M_{\mathrm{S}}-v_{\mathrm{k}}$ relation discussed in the previous subsection, the curves in the left panels are obtained assuming an SMBH of $M_{\mathrm{S}}=10^{6}$ M with a remnant of $v_{\mathrm{k}}=350$ km s-1, whereas curves in the right panels assume $M_{\mathrm{S}}=10^{7}$ M and $v_{\mathrm{k}}=250$ km s-1. We observe that, generally, $\Delta t_{\mathrm{cav}}$ is the dominant component, whereas the photon escape time $\Delta t_{\mathrm{ph}}$ is subdominant.
@@
Figure 8: Time lag of the counterpart flare after the GW event as a function of the radial location $a$ of the merger in the AGN disc. Curves in the left and right panels are obtained considering different SMBH accretion rates, whereas upper and lower panels depict solutions for different kick angles, as labelled. In each panel, the different curves correspond to flares in AGNs with different SMBH masses, as well as different magnitudes of remnant kick velocity, as indicated. The curves terminate/initiate when conditions 37 and $F_{\nu}/F_{\nu,\mathrm{host}}\leq 0.25$ are no longer satisfied (see the text for details).
Figure 9: Same as in Figure 8, but displaying the flare duration $\Delta t_{\mathrm{d}}$.